OpenStudy (anonymous):

match the parabolas with their focuses Focuses: (-2.75,-10), (2,11.75), (-4, -13.875), (1.25,17)

OpenStudy (anonymous):

not all of these will be used y=-x^2+4x+8 y=2x^2+16x+18 y=-2x^2+5x+14 y=-x^2+17x+7 y=2x^2+11x+5 y=-2x^2+6x+5

Mehek (mehek14):

what did you get so far?

OpenStudy (anonymous):

nothing, I don't even know how to begin

Mehek (mehek14):

use \(\dfrac{-b}{2a}\)

Mehek (mehek14):

\(ax^2+bx+c\)

Mehek (mehek14):

in the first equation, a = -1, b = 4

Mehek (mehek14):

plug that into the formula to get the x coordinate of the focus, then plug the value of x into the equation to find y

Mehek (mehek14):

what do you get for the first one?

OpenStudy (lynfran):

yea here an example for the 1st one the rest shd be done the same...